OpenStudy (pphalke):

A lab technician mixed a 660 ml solution of water and alcohol. if 3% of the solution is alcohol, how many mililiters of alcohol were used and how many mililiters of water were used?

OpenStudy (pphalke):

@jim_thompson5910 @fantageplayer @Tacolyfe

OpenStudy (pphalke):

@TheWiseLord

OpenStudy (decentnabeel):

A lab technician mixed a 660mL solution of water and alcohol Total solution: 660ml % of the solution is alcohol: 3% of 660 = 0.03 * 660 = 19.80ml how many mL of water total solution - amount of alkohol = 660-19.80 = 640.20ml

OpenStudy (decentnabeel):

@pphalke are you understand

OpenStudy (pphalke):

so 19.80 that would be the amount of alcohol and 640.20 that would be the amount of water right??

OpenStudy (pphalke):

@DecentNabeel